5/8 + 8/n + 3/n =
Solve for n
Solve for answer.
Thank you

Answer by josgarithmetic(39617)

This exercise problem is incomplete. The righthand member of the equation must have either a value or a variable, or there is not enough information to solve for n.
Can we assume that the left hand member is equal to some number, b? Then we may have 5/8 + 8/n + 3/n = b.

Let's clear the denominator by multiplying both sides by n.
n( 5/8 + 8/n + 3/n ) = bn
(5/8)n + 8 + 3 = bn
.... and can you finish this on your own, solving for n?